# Influencer Identification on Link Predicted Graphs

### 05 Feb 2024 v1

@@ Abstract
How would admissions look like in a university program for influencers? In the realm of social network analysis, influence maximization and link prediction stand out as pivotal challenges. Influence maximization focuses on identifying a set of key nodes to maximize information dissemination, while link prediction aims to foresee potential connections within the network. Given the complexity of these tasks, especially in large-scale networks, we propose a novel algorithm, The Social Sphere Model, which uniquely combines link prediction techniques with influence maximization strategies to effectively identify future vital nodes in weighted networks. Our approach is tested on two distinct contagion models, offering a promising solution with lower computational demands. This advancement not only enhances our understanding of network dynamics but also opens new avenues for efficient network management and influence strategy development.
@@ I Introduction
We seeks to fill a very important gap in the literature which would allow the understanding of influencer predictions in evolving social networks; that is, the effects of link prediction on heuristic centrality measures. To the best of our knowledge, link prediction similarity metrics have not yet been used with centrality metrics and $k$ node selection algorithms to approximate the effects of time on a dynamic social network.
@@ II.2 Link Prediction
A New Resource Allocation Variant: RA22-2- 2. Finally, returning to the intuition that common neighbors with larger degrees contribute less to the similarity between two nodes, we shall introduce here a new similarity metric, novel version of resource allocation. Suppose a node $u$ wants to connect two of its neighbors: there are then $\binom{|N(u)|}{2}$ pairs to choose from, so each pair of its neighbors has a $\frac{1}{\binom{|N(u)|}{2}}$ probability of being linked. For ease of calculation, we approximate this to
@@ II.4 Heuristic Top-$k$ Algorithms
VoteRank has been modified in the literature to fit weighted graphs by changing the score to include the importance of having many neighbors (e.g. see Sun et al. wvoterank ). Moreover, VoteRank has also been combined with metrics such as the DIL method (proposed by Liu et al. dil ) and coreness (e.g. see Li et al. dilvoterank and Kumar and Panda kcorevoterank ); inspired by these works, we give a generalization of VoteRank for a given centrality metric $c$ (such as the ones in Section II.3) by setting each node’s voting ability to its score with respect to $c$.
@@ III The Social Sphere Model
Expanding on the previous sections, we shall introduce here our novel model, the Social Sphere Model, an approach designed here to predict future influencers in changing social networks. This model synthesizes key elements of link prediction and top-$k$ influencer algorithms, offering a comprehensive tool for understanding and anticipating changes in social networks. The uniqueness of our approach lies in the integration of these two well-established methodologies to address the dynamic nature of social networks, which, to the best of our knowledge, has not been done before.
@@ III.2 Future Network Prediction
We shall combine these and look for future influencers by first predicting the state of the future network. This is achieved by calculating and normalizing link prediction similarity scores into probabilities of transmission, which are then assigned as weights for new edges. Moreover, given the unit of time $t$ we are predicting for, we can find a better predictor for the transmission probability.
@@ III.3 Future Top-$k$ Nodes Prediction
In what follows we shall first described the methods developed in order to introduce the Social Sphere Model and its main algorithm for future Top-$k$ Nodes Prediction, and then give an analysis of the results one may obtain through the model, as well as describe some of the applications and future directions.
@@ IV Methods
In order to illustrate the different uses of our model and our algorithm, we shall test our algorithm on both a random graph and an empirical dataset: given a graph $G$, we form a training graph from a random subset of existing edges of given size, predict the future state of the training graph, and identify influencers on those predicted graphs. For each graph and parameter tested, we repeated the process ten times, averaging the data. We modelled a dynamic network by taking the training graph as the current configuration and the original graph $G$ as the future network, thus allowing us to assess the efficacy of the influencers found in terms of their ability to “infect” nodes over time when set as the initial spreaders in the graph $G$.
@@ V.2 Results on an Empirical Dataset
Table 7: Average MSE over all top-$k$ algorithms for link prediction metrics using $70\%$ training graphs from the arXiv GrQc dataset cagrqc .
@@
Table 8: Average MSE over top-$k$ algorithms for link prediction metrics using $90\%$ training graphs from cagrqc .
@@
Mean Squared Error (MSE) Comparison. From Table 9 and Table 10, one can see that the MSE of the percentage infected is still small, though substantially larger than before, likely due to the larger variation that comes with more vertices and a larger graph. Notably, our new link prediction metric has the smallest average MSE over all the prediction metrics in both scenarios, with it also being the smallest in $75\%$ of the metrics in Table 10; A similar trend is in Table 11 and Table 12, where our new metric has the highest fraction in the $70\%$ case and the second highest in the $90\%$ scenario. These results support our hypotheses, as there is a significant decrease in MSE in the $90\%$ scenario as opposed to the $70\%$ one, this indicating that our link prediction metrics can identify influencers with very similar influence to that of the original, thus effectively showing prediction.
@@
The average performances of the various algorithms appears in Figure 6 and Figure 7. Notice that, as in the random graph case, VoteRank performs admirably in both simple and complex contagion in each scenario. Intriguingly, the $k$ highest algorithm also seems to do well, particularly in complex contagion. Unlike in Section V.1, LIR and LIR22-2- 2 exhibit sub-optimal efficiency and have effectively swapped with graph coloring and joint nomination, possibly because of the difference in structure of the empirical dataset and the Erdös-Rényi random graph. For instance, the empirical dataset is not fully connected (thus impossible for the fraction infected to be $1$, as there are too many disconnected pieces).
@@ VI Conclusion
In this research, we introduced the Social Sphere Model, a novel algorithm which utilizes link prediction and influencer identification to forecast future influencers in weighted networks. Our comprehensive analyses on both a random graph and a real world scenario (through the arXiv collaboration network) have demonstrated the algorithm’s efficacy, particularly evident in the low mean squared error and relatively high accuracy, suggesting the Social Sphere Model’s promising prediction capabilities for detecting future influencers. Moreover, our newly proposed similarity metric, RA-$2$, particularly excelled in both mean squared error and accuracy for the $90\%$ graphs. Moreover, our analysis of different link prediction metrics differing mostly in the degree of their denominators poses interesting questions around the optimal metric when applying the Social Sphere Model for different graphs and parameters.
The applications of our algorithm span a wide array of fields from marketing to epidemiology, offering valuable insights for strategic planning and information control:
  • Viral marketing: taking our algorithm over existing social media data and estimating $t$ from the expected time for production can assist in identifying potential future influencers in the network, giving more time (a valuable commodity) for negotiation, scouting, and strategizing (see Section I). The relatively short time complexities of heuristic models are also sometimes valuable in these scenarios. As an analogy, if one wanted to understand how admissions would look like in a university program for influencers, our algorithm would allow one to evaluate the potential of an influencer or group of influencers in the wider network.
@@
Future studies could also investigate a broader range of link prediction metrics beyond the neighborhood-based and quasi-local metrics employed here, like the clustering coefficient clusteringcoeff . Additionally, the concept of iteratively applying link prediction to evolving graphs, as in futurecommonneighbors , presents an intriguing avenue for extensions to our algorithm and potentially offers a more dynamic understanding of network changes over time. Our use of a variable $t$ in Algorithm 8 seeks to accommodate the increase in probability of interaction over time on the already predicted edges, but it does not account for any future edges that might have formed due to the predicted links’ existence. This paper also chose to include all potential edges, but limiting the additions to the highest-scoring pairs of vertices could increase precision by removing an excess of unlikely edges from posing issues for algorithms like the graph coloring method and the LIR methods.
Finally, it would be of much interest to study the implications of the Social Sphere Model on different network types: in particular, towards identifying the most effective combinations of link prediction and influencer identification metrics for specific network types, such as temporal or directed networks.
Acknowledgments. We extend our gratitude to the MIT PRIMES-USA program for their invaluable support and guidance throughout this research. The research of LPS is partially supported by NSF FRG Award DMS- 2152107, a Simons Fellowship, an NSF CAREER Award DMS 1749013 and a Visiting Fellowship at All Souls College, Oxford.
@@
    University of Illinois at Chicago, Chicago, USA.
@@
    Lakeside High School, Seattle, WA 98107, USA.

### 16 Feb 2024 v2

@@ Abstract
How would admissions look like in a university program for influencers? In the realm of social network analysis, influence maximization and link prediction stand out as pivotal challenges. Influence maximization focuses on identifying a set of key nodes to maximize information dissemination, while link prediction aims to foresee potential connections within the network. These strategies, primarily deep learning link prediction methods and greedy algorithms, have been previously used in tandem to identify future influencers. However, given the complexity of these tasks, especially in large-scale networks, we propose an algorithm, The Social Sphere Model, which uniquely utilizes expected value in its future graph prediction and combines specifically path-based link prediction metrics and heuristic influence maximization strategies to effectively identify future vital nodes in weighted networks. Our approach is tested on two distinct contagion models, offering a promising solution with lower computational demands. This advancement not only enhances our understanding of network dynamics but also opens new avenues for efficient network management and influence strategy development.
@@ I Introduction
In 2019, Ghafouri and Khasteh lpim1 introduced the idea of using a graph modelling technique to engage in link prediction prior to use of greedy algorithms for influence maximization as a way to account from invisible edges between nodes (i.e. edges not observed when collecting data on the network). Singh and Kailasam lpim2 utilized a similar approach when predicting sets of influential nodes for the next iteration of the graph through link prediction using the a type of Restricted Boltzmann Machine (RBM). In 2023, Yanchenko et al. exante discussed the use of deep learning link prediction to understand the outcomes of the future networks, then identify influential nodes through the greedy algorithm and dynamic degree discount. We take a similar approach, but focus on heuristic link prediction measures whereas they consider linear regression, deep learning, matrix factorization, and graph neural networks, and more. We also evaluate many other common centrality metrics and algorithms that, to the best of our knowledge, previous literature has not covered. In an effort to lower time complexity, our graph prediction also makes use of expected value to decrease computation cost of influence spread.
@@ II.2 Link Prediction
A New Resource Allocation Variant: RA22-2- 2. Finally, returning to the intuition that common neighbors with larger degrees contribute less to the similarity between two nodes, we shall introduce here a new similarity metric as a version of resource allocation. To the best of our knowledge, this is the first time this variant has been introduced. Suppose a node $u$ wants to connect two of its neighbors: there are then $\binom{|N(u)|}{2}$ pairs to choose from, so each pair of its neighbors has a $\frac{1}{\binom{|N(u)|}{2}}$ probability of being linked. For ease of calculation, we approximate this to
@@ II.4 Heuristic Top-$k$ Algorithms
Later on, researchers such as Sun et al. wvoterank have modified VoteRank to fit weighted graphs by changing the score to include the importance of having many neighbors. Researchers such as Li et al. dilvoterank and Kumar and Panda kcorevoterank have combined VoteRank with metrics such as the DIL method (proposed by Liu et al. dil ) and coreness; inspired by these works, we give a generalization of VoteRank for a given centrality metric $c$ (such as the ones in Section II.3) by setting each node’s voting ability to its score with respect to $c$.
@@ III The Social Sphere Model
Expanding on the previous sections, we shall introduce here our model, the Social Sphere Model for path-based similarity and heuristic centrality measures, an approach designed here to predict future influencers in changing social networks. This model synthesizes key elements of link prediction and top-$k$ influencer algorithms, offering a comprehensive tool for understanding and anticipating changes in social networks. The uniqueness of our approach lies in the use of lower complexity link prediction metrics rather than more costly neural networks and the inclusion of expected value in our predicted graph creation.
@@ III.2 Future Network Prediction
We shall combine these and look for future influencers by first predicting the state of the future network, like Yanchenko et al. exante did. However, when considering the formation of the future graph, rather than selecting the top edge pairs, we simply take normalized similarity scores to be probabilities of transmission and add them to the graph as weights, using expected value to calculate the distance. Moreover, given the unit of time $t$ we are predicting for, we can find a better predictor for the transmission probability.
@@ III.3 Future Top-$k$ Nodes Prediction
In what follows we shall first described the methods developed in order to implement our model, give an analysis of our results, and describe some of the applications and future directions.
@@ IV Methods
In order to illustrate the different uses of our model and our algorithm, we shall test our algorithm on both a random graph and an empirical dataset: given a graph $G$, we form a training graph from a random subset of existing edges of given size, predict the future state of the training graph, and identify influencers on those predicted graphs. For each graph and parameter tested, we repeated the process ten times, averaging the data. We modelled a dynamic network by taking the training graph as the current configuration and the original graph $G$ as the future network, thus allowing us to assess the efficacy of the influencers found in terms of their ability to “infect” nodes over time when set as the initial spreaders in $G$.
@@ V.2 Results on an Empirical Dataset
Table 7: Average MSE over all top-$k$ algorithms for link prediction metrics using $70\%$ training graphs from the arXiv GrQc dataset.
@@
Table 8: Average MSE over all top-$k$ algorithms for link prediction metrics using $90\%$ training graphs from the arXiv GrQc dataset.
@@
Mean Squared Error Comparison. Considering Table 9 and Table 10, one can see that the mean squared error of the percentage infected is still small, though substantially larger than before, likely due to the larger variation that comes with more vertices and a larger graph. Notably, our new link prediction metric has the smallest average MSE over all the prediction metrics in both scenarios, with it also being the smallest in $75\%$ of the metrics in Table 10; we also see a similar trend in Table 11 and Table 12, where our new metric has the highest fraction in the $70\%$ case and the second highest in the $90\%$ scenario. These results support our hypotheses, as there is a significant decrease in MSE in the $90\%$ scenario as opposed to the $70\%$ one, this indicating that our link prediction metrics can identify influencers with very similar influence to that of the original, thus effectively showing prediction.
@@
We have mapped out the average performances of the various algorithms in Figure 6 and Figure 7. Notice that, as in the random graph case, VoteRank performs admirably in both simple and complex contagion in each scenario. Intriguingly, the $k$ highest algorithm also seems to do well, particularly in complex contagion. Unlike in Section V.1, LIR and LIR22-2- 2 exhibit sub-optimal efficiency and have effectively swapped with graph coloring and joint nomination. A possible explanation for this observation is the difference in structure of the empirical dataset and the Erdös-Rényi random graph. For instance, the empirical dataset is not fully connected (and it is thus impossible for the fraction infected to be $1$, as there are too many disconnected pieces).
@@ VI Conclusion
In this research, we introduced the Social Sphere Model, a novel algorithm which utilizes heuristic link prediction and influencer identification to forecast future influencers in weighted networks. Our comprehensive analyses on both a random graph and a real world scenario (through the arXiv collaboration network) have demonstrated the algorithm’s efficacy, particularly evident in the low mean squared error and relatively high accuracy, suggesting the Social Sphere Model’s promising prediction capabilities for detecting future influencers, which is similar to Yanchenko et al. exante ’s findings that the influence of their model’s selected nodes were comparable to that of the Oracle method.
By modifying and extending several existing influencer identification algorithms to incorporate centrality metrics, we found that the modified forms of VoteRank were comparable to the original form of VoteRank when using metrics such as betweenness. Our proposed similarity metric, RA-$2$, particularly excelled in both mean squared error and accuracy for the $90\%$ graphs. Moreover, our analysis of different link prediction metrics differing mostly in the degree of their denominators poses interesting questions around the optimal metric when applying the Social Sphere Model for different graphs and parameters.
The practical applications of our algorithm span a wide array of fields from marketing to epidemiology, offering valuable insights for strategic planning and information control. These include:
  • Viral marketing: in the scenario mentioned in Section I, taking our algorithm over existing social media data and estimating $t$ from the expected time for production can assist in identifying potential future influencers in the network, giving more time (a valuable commodity) for negotiation, scouting, and strategizing. The relatively short time complexities of heuristic models are also sometimes valuable in these scenarios. As an analogy, if one wanted to understand how admissions would look like in a university program for influencers, our algorithm would allow one to evaluate the potential of an influencer or group of influencers in the wider network.
@@
In particular, our model differentiates from other ones by making use of relatively low time complexity metrics and algorithms as opposed to neural networks, the training of which may be infeasible for smaller businesses that might seek to calculate future influencers. The simplicity and efficiency of it can thus allow for quick identification of influencers and thus better networking.
@@
Finally, it would be of much interest to study the implications of the Social Sphere Model more graph theoretically on different network types: in particular, towards identifying the most effective combinations of link prediction and influencer identification metrics for specific network types, such as temporal or directed networks, varying empirical dataset origins, and properties like varying clustering coefficient or diameter values.
Acknowledgments. We extend our gratitude to the MIT PRIMES-USA program for their invaluable support and guidance throughout this research. The research of LPS is partially supported by NSF FRG Award DMS- 2152107, a Simons Fellowship and NSF CAREER Award DMS 1749013. Part of the work appearing here was done whilst LPS was a Visiting Fellow at All Souls College, Oxford.
@@
    University of Illinois at Chicago, Chicago, IL 60607, USA.
@@
    Lakeside High School, Seattle, WA 98107.